OpenStudy (anonymous):

A trapezoid has an area of 22 square inches. The longer base measures 6 inches and the height is 4 inches. use the formula A=1/2h(b1+b2) to find the length of the shorter base.

OpenStudy (anonymous):

oh sheet I hate these.

OpenStudy (anonymous):

oh, actually that was easy b2=5

OpenStudy (anonymous):

22=1/2(4)(6+b2) 22=2(6+b2) 22=12+2b2 (Distribute) 10=2b2 b2=5

OpenStudy (anonymous):

lol i know i hate these ones too.thank you(:

OpenStudy (anonymous):

np